OpenStudy (anonymous):

the length of a rectangular pizza is 4 in. less that twice its width. The area of the pizza is 160 in.^2. Find the dimensions of the pizza

OpenStudy (anonymous):

width=w length=2w-4 area=length*width 160=(2w-4)(w) 2w^2-4w-160=0 w^2-2w-80=0 (w-10)(w+8)=0 w=10, -8 length can't be negative so w=10, l=16